Studies show that massage therapy has a variety of health benefits and it is not too expensive. A sample of 12 typical one-hour massage therapy sessions showed an average charge of $60. The population standard deviation for a one-hour session is σ =$5.1

a) Using 90% confidence, what is the margin of error (to 2 decimals)?

b) Using 99% confidence, what is the margin of error (to 2 decimals)?
Answer #1

Solution:-

a) The margin of error for 90% confidence is 2.42.

M.E = z_{\alpha /2}\times{\frac{\sigma }{\sqrt{n}}}

M.E = 1.645*1.47224

M.E = 2.422

b) The margin of error for 99% confidence is 3.80.

M.E = z_{\alpha /2}\times{\frac{\sigma }{\sqrt{n}}}

M.E = 2.576*1.47224

M.E = 3.7925